Derive the value of a3+b3 from ( a+b)3 and what is the relation between (a+b)2 and( a-b) means find the solution of (a+b)2 in form( a-b)


1
Expert's answer
2021-07-12T16:42:16-0400

First of all let us know what is "(a+b)^3"


"(a+b)^3= (a+b) (a+b) (a+b)"


"={(a+b) (a+b)} (a+b)"

"={a(a+b) + b(a+b)} (a+b)"


"=(a^2 + ab + ab + b^2) (a+b)"


"=(a^2 + b^2 + 2ab) (a+b)"


"=a^2(a+b) + b^2(a+b) + 2ab(a+b)"


"=a^3 + a^2b + ab^2 + b^3 + 2a^2b + 2ab^2"


"=a^3 + b^3 + 3a^2b + 3ab^2"


"=a^3 + b^3 + 3ab(a+b)"


Now when we have expanded  "(a+b)^3 = a^3 + b^3 + 3ab(a+b)"


We can equate it


"(a+b)^3 = a^3 + b^3 + 3ab(a+b)"


"(a+b)^3 - 3ab(a+b) = a^3 + b^3"


"\therefore a^3 + b^3 = (a+b)^3 - 3ab(a+b)"


The following is the relation between "(a+b)^2" and "( a-b)"


"(a+b)^2=(a-b)+a(a-1)+b(2a+b+1)"


 The solution of "(a+b)2" in form "( a-b)"


"(a+b)^2=(a+b)(a+b)"


"=a(a+b)+b(a+b)"


"=a^2+ab+ab+b^2+a-a+b-b"


Rearranging gives


"(a+b)^2=(a-b)+a^2-a+2ab+b^2+b"


Factor common terms

"\therefore (a+b)^2=(a-b)+a(a-1)+b(2a+b+1)"